OpenStudy (anonymous):

Solve for x: 3 over 4 x + 5 over 8 = 4x

OpenStudy (anonymous):

plz help

OpenStudy (compassionate):

\[\frac{ 3 }{ 4x } + \frac{ 5 }{8 } = \frac{ 1 }{ 4x }\] Cross multiply.

OpenStudy (anonymous):

thank you ill tell u if i passed the quiz

OpenStudy (anonymous):

yaaay i passed

OpenStudy (compassionate):

Congrats!